stap dows 2 Ve ½Ta fnd rad circut는.tat.., o.c

i need only wrie this page in computer font
Answer #1

Transformer is a magnetic device which couples two or more winding.Transformer AC operates AC excitation and change voltage level when side to other side.

N1>N2 Step down

N1<N2 Step up

\frac{V1}{V2}=\frac{N1}{N2}=\frac{I2}{I1}

Short circuit test:

Z=\frac{Vsc}{Ioc}\Omega

X=(Z^2-R^2)^\frac{1}{2}

R=\frac{Psc}{Isc^2^{}}

Open circuit test:Zc=\frac{Poc}{Voc^2}

\frac{1}{Z}=Y=\frac{Ioc}{Voc}

bm=(Y^2-\sigma ^2)^\frac{1}{2}

12 Primary Secondary